OpenStudy (xapproachesinfinity):

Hey! I'm thinking of deriving 1+4+9+....+n^2=[n(n+1)(2n+1)]/6 directly no imduction i was thinking of the series \(X_{n+1}-X_n\) which give me 3+5+7+9+11+13+......+k which is the sum of the first n odd number that sums to k^2-1 how will this help. Give me just a hint lol

OpenStudy (xapproachesinfinity):

what last term should be 2k-1 not k lol for the sum of the first odd intergers

OpenStudy (xapproachesinfinity):

that*

ganeshie8 (ganeshie8):

that gives you a formula for sum of first n natural numbers you can use that to find sum of squares of first n natural numbers

ganeshie8 (ganeshie8):

\[\large \sum \limits_{k=1}^n (k+1)^2 - k^2 = (n+1)^2 - 1\]

OpenStudy (xapproachesinfinity):

the formula for n natural numbers is n(n+1)/2 the one i was thinking about is sum of n first positive integers

ganeshie8 (ganeshie8):

\[\large \sum \limits_{k=1}^n (2k+1) = (n+1)^2 - 1\] \[\large \sum \limits_{k=1}^n 2k+ \sum \limits_{k=1}^n 1 = (n+1)^2 - 1\] \[\large \sum \limits_{k=1}^n 2k+ n = (n+1)^2 - 1\] simplifying a bit you get \[\large \sum \limits_{k=1}^n k = \dfrac{n(n+1)}{2}\]

ganeshie8 (ganeshie8):

use the same trick to find the sum of squares consider \(\sum\limits_{k=1}^n (k+1)^3 - k^3\) to start with

ganeshie8 (ganeshie8):

telescoping gives you \[\sum\limits_{k=1}^n \left((k+1)^3 - k^3\right) = (n+1)^3-1 \]

ganeshie8 (ganeshie8):

\[\sum\limits_{k=1}^n \left(k^3+3k^2+3k+1 - k^3\right) = (n+1)^3-1 \]

ganeshie8 (ganeshie8):

\[\sum\limits_{k=1}^n \left(3k^2+3k+1\right) = (n+1)^3-1 \]

ganeshie8 (ganeshie8):

see if u can take it home

OpenStudy (xapproachesinfinity):

eh so what is started with sum of (Xn+1-Xn) which generated the the sum of first n odd integer is what you used! to find the sum of first n natural number I thought that this would work for sum of the squares

ganeshie8 (ganeshie8):

Exactly!

ganeshie8 (ganeshie8):

as you can you can use the same trick to find sum of ANY powers of first n natural numbers

OpenStudy (xapproachesinfinity):

eh i guess from i get it! let me see

ganeshie8 (ganeshie8):

the only drawback of this method is that you need to know formulas for ALL the lowest powers

OpenStudy (xapproachesinfinity):

That's it! I got I already have the sum of k and the sum of (1) I just need to factor the RHD and push the term of the left to the right leaving sum of k^2

ganeshie8 (ganeshie8):

Yep! see if u can find a formula for below if u have time : \[1^4 + 2^4 + 3^4 + \cdots + n^4 = \dfrac{n(n+1)(2n+1)(3n^2+3n-1)}{30}\]

OpenStudy (xapproachesinfinity):

this part was not easy to come up with though lol \(\sum\limits_{k=1}^n \left((k+1)^3 - k^3\right) = (n+1)^3-1\)

ganeshie8 (ganeshie8):

actually i think thats the only interesting part here

ganeshie8 (ganeshie8):

expand that sum and you will see lot of cancellation..

ganeshie8 (ganeshie8):

\[\sum\limits_{k=1}^n \left((k+1)^3 - k^3\right) = (2^3 - 1^3) +(3^3-2^3) + (4^3-3^3) + \cdots + (n+1)^3 - n^3\]

OpenStudy (xapproachesinfinity):

oh yeah there is cancellations interesting

OpenStudy (xapproachesinfinity):

ok for the other one i have to start with \(\sum (k+1)^4-k^4=(n+1)^4-1\)

ganeshie8 (ganeshie8):

In general \[\sum\limits_{k=1}^n \left(f(r+1) - f(r)\right) = f(r+1) - f(1)\] this sum is so interesting that it has a name, its called `telescoping sum`

OpenStudy (xapproachesinfinity):

i just realized that i was approaching the same thing with Xn+1-Xn taking the sum i just needed to telescope things

ganeshie8 (ganeshie8):

yes thats a good start, also u need to find sums of cubes upfront.. :)

OpenStudy (xapproachesinfinity):

oh ok! i get the general idea which what I'm looking thanks alot

OpenStudy (zarkon):

\[\sum\limits_{k=1}^n \left((k+1)^3 - (k-1)^3\right) = (n+1)^3+n^3-1\] \[\sum\limits_{k=1}^n \left((k+1)^3 - (k-1)^3\right) =\sum\limits_{k=1}^n \left(6k^2+2\right)\] \[6\sum\limits_{k=1}^n k^2+2n= (n+1)^3+n^3-1\] \[6\sum\limits_{k=1}^n k^2= (n+1)^3+n^3-1-2n=n(n+1)(2n+1)\]

OpenStudy (zarkon):

then you don't need to know \[\sum\limits_{k=1}^nk=\frac{n(n+1)}{2}\]

OpenStudy (xapproachesinfinity):

cool^_^ series are really interesting hehe

ganeshie8 (ganeshie8):

thats a clever trick to avoid computing \(\sum k\) xD

OpenStudy (xapproachesinfinity):

\(\sum_{i=1}^{ n} (k+1)^4-k^4=(n+1)^4-1\) this one i wrote above is giving me the sum of k^3

ganeshie8 (ganeshie8):

Oh yes because k^4's cancel out on left hand side

OpenStudy (xapproachesinfinity):

right^_^

OpenStudy (xapproachesinfinity):

i suppose for k^4 I should do sum of (k+1)^5-k^5 if the pattern continues hehehe

ganeshie8 (ganeshie8):

yes:)

OpenStudy (xapproachesinfinity):

Cool! thanks ^_^

ganeshie8 (ganeshie8):

Also notice, using @Zarkon 's trick, you can avoid computing all previous odd powers

ganeshie8 (ganeshie8):

\[\sum\limits_{k=1}^n \left((k+1)^5 - (k-1)^5\right) = (n+1)^5+n^5-1\] \[\sum\limits_{k=1}^n \left(10k^4+20k^2+2\right) = (n+1)^5+n^5-1\]

OpenStudy (xapproachesinfinity):

Eh! i thought that won't work for every case lol

ganeshie8 (ganeshie8):

it works always... we can convince it ourselves by playing with sums ..

ganeshie8 (ganeshie8):

maybe lets do a quick derivation

OpenStudy (xapproachesinfinity):

hmm ok what derivation

ganeshie8 (ganeshie8):

prove \[\sum\limits_{k=1}^n \left((k+1)^5 - (k-1)^5\right) = (n+1)^5+n^5-1 \]

OpenStudy (xapproachesinfinity):

Inductions i guess lol

ganeshie8 (ganeshie8):

its a one line proof, trust me

OpenStudy (xapproachesinfinity):

i have to go sleep lol i will work on this tomorrow! =========== i was thinking of adding that n^5 to both side and manipulate it in the LHS

OpenStudy (xapproachesinfinity):

taking into count the what we did before of course!

ganeshie8 (ganeshie8):

\[\begin{align}\sum\limits_{k=1}^n \left((k+1)^5 - (k-1)^5\right) &= \sum\limits_{k=1}^n \left((k+1)^5 - k^5 + k^5- (k-1)^5\right) \\~\\ &= \sum\limits_{k=1}^n \left((k+1)^5 - k^5\right) + \sum\limits_{k=1}^n \left( k^5- (k-1)^5\right) \\~\\ \end{align}\]

OpenStudy (xapproachesinfinity):

i was trying to do the opposite hehe

ganeshie8 (ganeshie8):

\[\begin{align}\sum\limits_{k=1}^n \left((k+1)^5 - (k-1)^5\right) &= \sum\limits_{k=1}^n \left((k+1)^5 - k^5 + k^5- (k-1)^5\right) \\~\\ &= \sum\limits_{k=1}^n \left((k+1)^5 - k^5\right) + \sum\limits_{k=1}^n \left( k^5- (k-1)^5\right) \\~\\ &= \left((n+1) ^5-1\right) +\left( n^5-0\right) \end{align}\]

ganeshie8 (ganeshie8):

yeah anything works! have good sleep :)

OpenStudy (xapproachesinfinity):

eh i tried to go the other way, I was almost close alright thanks for the help and thanks everyone^_^